# Continual Learning: Forget-free Winning Subnetworks for Video Representations

### 19 Dec 2023 v1

@@ 6.2 WSN+FSO’s Video Representations
Figure 9: The comparisons of WSN of NeRV Blocks (blue area) with FSO (white area) in terms of Feature variances and high-frequency components: the (a) offers the variance of the feature map and the (b) provides $\Delta\log$ amplitudes at high-frequency ($1.0\pi$).
@@ 7 Results On Few-shot CIL (FSCIL)
7.1 Results and Comparisons
Table VI: Classification accuracy of ResNet18 on CIFAR-100 for 5-way 5-shot incremental learning with the same class split as in TOPIC [14]. ${}^{\ast}$ denotes the results reported from [71]. ${}^{\dagger}$ represents our reproduced results.
@@ 7.1 Results and Comparisons
Table VIII: Classification accuracy of ResNet18 on CUB-200-2011 for 10-way 5-shot incremental learning (TOPIC class split [81]). ${}^{\ast}$ denotes results reported from [71]. ${}^{\dagger}$ represents our reproduced results.

### 04 Jan 2024 v2

@@ 6.2 WSN+FSO’s Video Representations
Figure 9: Comparisons of WSN of NeRVs (blue area) with FSO (white area) in terms of Feature variances and high-frequency components: the (a) offers the variance of the feature map and the (b) provides $\Delta\log$ amplitudes at high-frequency ($1.0\pi$).
@@ 7 Results On Few-shot CIL (FSCIL)
Table VI: Classification accuracy of ResNet18 on CIFAR-100 for 5-way 5-shot incremental learning with the same class split as in TOPIC [14]. ${}^{\ast}$ denotes the results reported from [72].
@@
Table VIII: Classification accuracy of ResNet18 on CUB-200-2011 for 10-way 5-shot incremental learning (TOPIC class split [82]). ${}^{\ast}$ denotes results reported from [72].